# Free Electron Paths from Dirac’s Wave Equation Elucidating Zitterbewegung and Spin

### v2

@@ 1 Introduction
The Dirac relativistic wave equation (Dirac (1928)) is arguably the best particle model available for the dynamics of a single electron in the absence of particle creation and annihilation. It correctly predicts that the electron has spin with magnitude $\hbar/2$ and it has a magnetic moment with $g=2$ that is very close to the experimentally observed value, although it does not seem to reveal the physics behind these values. It also accurately predicts the electro-magnetic spectrum for the hydrogen atom (Dirac (1928)). However, when Schrödinger performed a Heisenberg operator analysis of Dirac’s equation for a free electron, he found that it exhibits a component of motion with an ultra-high frequency $\omega_{0}=2mc^{2}/\hbar\sim 1.55\times 10^{21}$s-1 that he called zitterbewegung (trembling motion) (Schrödinger (1930)). Various explanations have been proposed over the last century for the source of this puzzling phenomenon but none have been widely accepted. A difficulty associated with it is that the electron appears to move at the speed of light $c$, even though it has a non-zero mass $m$; in fact, Dirac’s velocity operators that appear in his equation have eigenvalues of magnitude $c$. In this work, we will show that, contrary to common wisdom, Dirac’s equation can reveal the physics behind these puzzling features.
@@
Although these classically-based models have been available for a long time, they have not had much impact on mainstream physics where Schrö-dinger’s zitterbewegung, if it is mentioned at all, is usually described as ”the interference between positive and negative energy solutions of Dirac’s equation”, which is based on the mathematics that produces the zitter motion but which lacks a physical picture of what is behind it. Actually, more than 70 years ago, Huang (1952) did provide a physical explanation of zitterbewegung based on his analysis of a wave function expressed as a superposition of “positive energy” and “negative energy” solutions of Dirac’s wave equation for a free electron. His analysis agrees with the predicted behavior of the electron given by the afore-mentioned classical Dirac particle models. Bohm and Hiley (1993) give a similar interpretation of Pauli’s wave equation solution for the free electron in Chapter 10.4 of their book. They conclude that the spin angular momentum and magnetic moment of an electron come from a circulation of the electron as a point particle. These earlier studies have also had limited impact on the understanding of spin and zitterbewegung in mainstream physics.
@@
In this work, we show that, despite the widespread belief that Dirac’s equation does not reveal electron paths, they are hiding in plain sight. We derive the electron’s worldline from the Dirac wave function for a free electron, that is, its space-time path as a function of a proper time, and show that the nature of zitterbewegung and spin are exactly as predicted by the afore-mentioned classical Dirac particle models. In fact, we show that the model of Barut and Zanghi (1984), which is based on a Lagrangian function involving both “classical” spinors and classical dynamic variables, is equivalent to Dirac’s equation in the case of a free electron, that is, every Barut-Zanghi spinor solution is a Dirac spinor solution, and conversely.
@@
It has already been shown in Beck (2023) that there is an equivalence between the equations of motion of his neoclassical model of the electron and the equations of motion in classical dynamic variables that are derived by applying Dirac’s velocity and spin operators to the Barut-Zanghi spinor. Furthermore, as noted by Rodriguez et al. (1993), the Barut and Zanghi (1984) model is equivalent to the light-like zitter model of Hestenes (1993, 2010), which also uses spinors with classical dynamical variables to express the dynamics of an electron. Beck’s neoclassical model has the same Lagrangian function, given later, as the classical Dirac model of Salesi (2002) when minimal coupling is added to the latter. It also gives a covariant form of the equations of motion for the relativistic spinning particle model of Rivas (2003) that he derived from the Frenet-Serret equations for a curve in three-dimensional space. Therefore, all these classical Dirac particle models are equivalent to each other in their description of the dynamics of the spinning electron, and for the free electron, they all agree in this sense with Dirac’s wave function model. Furthermore, the electron’s motion in the classical Dirac models can be decomposed into a global part, which is the motion of the inertia (or spin) center, and a local spin part that is Schrödinger’s zitterbewegung, and we show that for the electron’s velocity, this corresponds to the Gordon decomposition of the Dirac current.
@@ 2 Solution of Covariant Dirac Equation for a Free Electron
using the normalization of $A$ given above and Schrödinger’s zitterbewegung frequency $\omega_{0}=2mc^{2}/\hbar$. This equation can be interpreted as the velocity of the electron when it is at the space-time point $x=(ct,\mathbf{x})$.
@@
As noted by Salesi (2002) but not proved by him, Eq. (2.9) corresponds to the covariant form of the Gordon decomposition of the Dirac current (dropping the charge $q$ as a multiplier)
@@ 2.1 Electron’s Wave Function and Space-time Motion in Proper Time
where $\omega_{1}=\omega_{0}/2$, showing that $\phi(0)=A$. In fact, this is just the wave function $\psi_{c}(\tau)$ expressed relative to $X_{c}$ boosted to give the wave function $\psi(x)=\phi(\tau(x))$ relative to $X_{o}$. Deboosting from $X_{o}$ to $X_{c}$ transforms $x$ to $\tau$ because Eq. (2.12) corresponds to the time part of the Lorentz transformation of $x=(ct,\mathbf{x})$ to $x_{c}=(c\tau,\mathbf{r})$, as shown in Beck (2023) in his Section 2.2. There is no dependence of the rest-frame wave function on $\mathbf{r}$ because $\mathbf{P}_{c}=\mathbf{0}$.
@@
The 4-vector proper velocity $u^{\mu}(\tau)$ is therefore a superposition of the constant global velocity $v^{\mu}=\pi^{\mu}/m$ and an oscillatory velocity $w^{\mu}(\tau)$ with angular frequency $\omega_{0}$ that is a manifestation of Schrödinger’s zitterbewegung. The same expression for the proper velocity was derived by Barut and Zanghi (1984), Salesi (2002) and Beck (2023) for their respective classical models of the Dirac electron. Notice that with the new notation, the previous velocity fields $\widetilde{u}^{\mu}(x)$ and $\widetilde{w}^{\mu}(x)$ can be written as $u^{\mu}(\tau(x))$ and $w^{\mu}(\tau(x))$; and similarly, $\psi(x)$ is $\phi(\tau(x))$.
@@
where the constant $y^{\mu}(0)=x^{\mu}(0)+\dot{u}^{\mu}(0)/\omega_{0}^{2}=x^{\mu}(0)-z^{\mu}(0)$. Actually, this interpretation is consistent with the Heisenberg operator for rate of change with respect to proper time that was introduced by Beck (1942) and also derived rigorously by Barut and Thacker (1985). Here, the Beck derivative operator gives
@@
and so $\dot{x}^{\mu}(\tau)=\bar{\phi}(\tau)~\hat{\dot{x}}^{\mu}\phi(\tau)=\bar{\phi}(\tau)~\hat{u}^{\mu}\phi(\tau)=u^{\mu}(\tau)$.
@@ 2.3 Rest-frame Zitter Motion of Electron
We choose origin $C$ to be the point $\mathbf{y}(\tau)$ relative to $X_{o}$ so that it moves with the global motion of the electron relative to $X_{o}$. This means that $C$ is chosen as the point where the momentum of the electron is located ($\mathbf{P}=m\mathbf{v}=m\dot{\mathbf{y}}$, not $m\mathbf{u}$). Clearly, the inertia of the electron is located at the point $C$ and not at the electron’s spatial position $\mathbf{x}(\tau)$ where the charge is located, so $C$ is appropriately called the inertia center (or center of mass) of the electron, although, as we will see, it is also appropriate to call it the spin center. This physical separation of the inertia and charge centers of an electron is a characteristic of the afore-mentioned classical Dirac particle models (Rivas (2015) focuses on this feature for his model). Furthermore, since the free electron is perpetually accelerating (see Eq. (2.15)), it is clear that the electron does not emit or absorb radiation (photons) unless its inertia center accelerates.
@@
This is consistent with $\phi^{*}(\tau)\phi(\tau)=\frac{1}{c}\bar{\phi}(\tau)\hat{u}^{0}\phi(\tau)=\frac{1}{c}u^{0}(\tau)=\dot{t}(\tau)=1$ because $t=\tau$ in the rest-frame at $C$.
@@
It is readily shown from Eq. (2.29) that $\mathbf{u}\cdot\mathbf{n}=0$ so that the electron’s zitter motion relative to the rest-frame $X_{c}$ lies in a plane orthogonal to the spin direction $\mathbf{n}$. Furthermore, $\mathbf{u}\cdot\mathbf{u}=c^{2}$ so the electron moves at the speed of light, and $\dot{\mathbf{u}}\cdot\dot{\mathbf{u}}=c^{2}\omega_{0}^{2}$, which justify the constraints $C1:u^{\mu}u_{\mu}=0$ and $C2:\dot{u}^{\mu}\dot{u}_{\mu}=-c^{2}\omega_{0}^{2}$ that Beck (2023) imposed on his neoclassical model. It is shown in Appendix A that $C1$ implies that $\mathbf{r}\cdot\mathbf{r}$ is constant. Thus, the zitter motion is a circle about $C$, which clearly has circular frequency $\omega_{0}$ and therefore radius $r_{0}=c/\omega_{0}=\hbar/(2mc)$.
This circular zitter motion produces a spin $\mathbf{s}$ of magnitude $\hbar/2$ in the direction $\mathbf{n}$ orthogonal to the plane of the motion. Furthermore, if the electron is at space-time point $x$, Eq. (2.20) shows that its angular position in this circular motion is $\omega_{0}\tau(x)=2\theta(x)$, using Eq. (2.12). Thus, as noted by Hestenes (1993), the electron’s angular position in its zitter motion is twice its wave function phase $\theta(x)$; for example, if this phase increases by $\pi$, then the electron will complete a full cycle around its circular zitter path.
@@ 3 Equivalence for Free Electrons of the Dirac Equation and the Classical Dirac Particle Models
This evolution equation is satisfied by the “classical spinor” in the electron model of Barut and Zanghi (1984), which implies that for a free electron, the Dirac wave function written as a function of proper time is a Barut-Zanghi spinor. Beck (2023) has proved the converse: given a Barut-Zanghi spinor $\phi(\tau)$, the spinor $\psi(x)=\phi(\tau(x))$, where $\tau(x)$ is defined in Eq. (2.12), satisfies Dirac’s wave equation and so is a Dirac wave function. This is obviously true here from reversing the procedure to get Eq. (2.13), that is, $\phi(\tau)$ in this equation is the spinor solution of Eq. (3.1), and so from Eq. (2.12), $\phi(\tau(x))=\psi(x)$, the solution of Dirac‘s equation given in Eq. (2.2).
@@
The free electron models of Hestenes (2010) and Barut and Zanghi (1984) are also equivalent, as noted by Rodriguez et al. (1993), so the same equivalence with Dirac’s wave equation also applies to Hestenes zitter (light-like) model. Both of these models use a mix of “classical spinors” and classical dynamic variables in their Lagrangian functions. The difference is that Hestenes uses his Space-time Algebra, a Clifford algebra, to express and analyze the Lagrangian function for his zitter model. Note that when there is no em-field, the Pauli dipole term that Hestenes adds to his Lagrangian plays no role in the dynamics.
Barut and Cruz (1993) also examine the addition of this dipole term to the Barut-Zanghi Lagrangian function with the goal of empirically producing the “anomalous” magnetic moment that is observed experimentally, which corresponds to a gyromagnetic ratio with $g\approx 2(1+\alpha/(2\pi))$ where $\alpha$ is the usual fine structure constant. It is important to note, however, that Dirac’s equation does not have an explicit dipole term since it uses only minimal coupling to the em-field but it still gives rise to magnetic and electric dipole moments, albeit with a gyromagnetic ratio corresponding to exactly $g=2$.
@@
giving two coupled second-order equations of motion describing the free electron that are derived from Dirac‘s equation. The equations of motion in Eq. (3.2) are a special case of the neoclassical model of Beck (2023) where Newton’s second law in proper time gives the global acceleration $\ddot{y}^{\mu}(\tau)$, consistent with the inertia being located at the inertia center $C$, the point whose spatial position is given by $\mathbf{y}(\tau)$ relative to the inertial reference frame $X_{o}$. His fundamental equations of motion are
@@
where $A=\left(V/c,\mathbf{A}\right)$ is the 4-vector potential for the em-field. This Lagrangian function is just the classical relativistic one with minimal coupling to the em-field (e.g. Goldstein (1959)) except for the additional last term, which turns out to be the contribution from the spin kinetic energy. The same Lagrangian was chosen for a free electron by Riewe (1972) and also by Salesi (2002) for his classical Dirac particle model. We refer to these electron models that are based on this Lagrangian as neoclassical relativistic mechanics models because they use classical relativistic mechanics with the addition of a new spin energy term in the Lagrangian function.
Rivas (2003) has presented a similar relativistic model of the electron that uses an inertial observer’s time $t$ rather than the proper time $\tau$. He derives it from the Frenet-Serret equations for a curve in three-dimensional space. By examining his equations of motion for a free electron relative to the rest-frame $X_{c}$, which is inertial, it can be shown that they are equivalent to the space part of Eq. (3.3) from the neoclassical model of Beck (2023). The time part in Eq. (3.3) is not needed in the rest-frame because observer time $t$ equals proper time $\tau$ there. For the interaction with an em-field, both of these theories use Newton’s second law for the motion of the inertia center, with the Lorentz force evaluated at the charge center. Therefore, the neoclassical model of Beck can be viewed as a covariant form of the relativistic model of Rivas.
Beck (2023) introduced a physically-based spin tensor $S$ defined by
@@
These same equations were derived by Barut and Zanghi (1984) by applying Dirac’s 4-vector velocity operator and his spin tensor operator to their “classical spinor” $\phi(\tau)$ (they did not give any proof, which is given in Section 4.2 of Beck (2023)). Thus, the simple electron model given in Eq. (3.3) has the same dynamics as the Barut-Zanghi model, and so for a free electron, the same dynamics as Dirac’s electron model. Not only are the equations in Eq. (3.7) implied by Dirac’s equation and operators for a free electron, they also hold for an electron in an em-field provided the dynamic variables $u^{\mu},S^{\mu\nu}$ and $\pi^{\mu}$ are replaced by their corresponding Dirac operators. This can be shown by employing the proper time derivative operator introduced by Beck (1942) to derive the Heisenberg operator equations from Dirac’s equation (e.g. see Barut (1987)).

### v3

@@ 1 Introduction
The Dirac relativistic wave equation (Dirac (1928)) is arguably the best particle model available for the dynamics of a single electron in the absence of particle creation and annihilation. It correctly predicts that the electron has spin with magnitude $\hbar/2$ and it has a magnetic moment with $g=2$ that is very close to the experimentally observed value, although it does not seem to reveal the physics behind these values. It also accurately predicts the electro-magnetic spectrum for the hydrogen atom (Dirac (1928)). However, when Schrödinger performed a Heisenberg operator analysis of Dirac’s equation for a free electron, he found that it exhibits a component of motion with an ultra-high frequency $\omega_{0}=2mc^{2}/\hbar\sim 1.55\times 10^{21}$s-1 that he called zitterbewegung (trembling motion) (Schrödinger (1930)). Various explanations have been proposed over the last century for the source of this puzzling phenomenon but none have been widely accepted. A difficulty associated with it is that the electron appears to move at the speed of light $c$, even though it has a non-zero mass $m$; in fact, Dirac’s velocity operators that appear in his equation have eigenvalues of magnitude $c$. In this work, we will show that, contrary to common wisdom, Dirac’s equation can reveal the physics behind these puzzling features. This analysis gives a resolution of several mysteries; for example, what is Schrödinger’s zitterbewegung that appears to have the electron moving at the speed of light; what is spin if it is not a rotation of the electron itself; and why does the electron, an apparent point charge, have dipoles?
@@
Although these classically-based models have been available for a long time, they have not had much impact on mainstream physics where Schrö-dinger’s zitterbewegung, if it is mentioned at all, is usually described as ”the interference between positive and negative energy solutions of Dirac’s equation”, which is based on the mathematics that produces the zitter motion but which lacks a physical picture of what is behind it. Actually, more than 70 years ago, Huang did provide a physical explanation of zitterbewegung based on analyzing a wave function expressed as a superposition of “positive energy” and “negative energy” solutions of Dirac’s wave equation for a free electron (Huang (1952)). His analysis agrees with the predicted behavior of the electron given by the afore-mentioned classical Dirac particle models. Bohm and Hiley give a similar interpretation of Pauli’s wave equation solution for the free electron in Chapter 10.4 of their book (Bohm and Hiley (1993)). They conclude that the spin angular momentum and magnetic moment of an electron come from a circulation of the electron as a point particle. These earlier studies have also had limited impact on the understanding of spin and zitterbewegung in mainstream physics.
@@
In this work, we show that, despite the widespread belief that Dirac’s equation does not reveal electron paths, they are hiding in plain sight. We derive the electron’s worldline from the Dirac wave function for a free electron, that is, its space-time path as a function of a proper time, and show that the nature of zitterbewegung and spin are exactly as predicted by the afore-mentioned classical Dirac particle models. In fact, we show that the model in Barut and Zanghi (1984), which is based on a Lagrangian function involving both “classical” spinors and classical dynamic variables, is equivalent to Dirac’s equation in the case of a free electron, that is, every Barut-Zanghi spinor solution is a Dirac spinor solution, and conversely.
@@
It has already been shown in Beck (2023) that there is an equivalence between the equations of motion of Beck’s neoclassical relativistic mechanics model of the electron and the equations of motion in classical dynamic variables that are derived by applying Dirac’s velocity and spin operators to the Barut-Zanghi spinor. Furthermore, as noted in Rodriguez et al. (1993), the Barut and Zanghi model is equivalent to Hestenes’ light-like zitter model in Hestenes (1993, 2010), which also uses spinors with classical dynamical variables to express the dynamics of an electron. Beck’s neoclassical model has the same Lagrangian function, given later, as the classical Dirac model in Salesi (2002) when minimal coupling is added to the latter. It also gives a covariant form of the equations of motion for the relativistic spinning particle model of Rivas that he derived from the Frenet-Serret equations for a curve in three-dimensional space (Rivas (2003)). Therefore, all these classical Dirac particle models are equivalent to each other in their description of the dynamics of the spinning electron, and for the free electron, they all agree in this sense with Dirac’s wave function model. Furthermore, the electron’s motion in the classical Dirac models can be decomposed into a global part, which is the motion of the inertia (or spin) center, and a local spin part that is Schrödinger’s zitterbewegung, and we show that for the electron’s velocity, this corresponds to the Gordon decomposition of the Dirac current.
@@ 2 Solution of Covariant Dirac Equation for a Free Electron
using the normalization of $A$ given above and Schrödinger’s zitterbewegung frequency $\omega_{0}=2mc^{2}/\hbar$. This equation can be interpreted as the velocity of the electron when it is at the space-time point $x=(ct,\mathbf{x})$. This interpretation is not novel but it is not the standard interpretation which regards $\widetilde{u}^{\mu}(x)$ as the expected value of the velocity if the electron is located at $x$.
@@
As noted in Salesi (2002) but not proved there, Eq. (2.9) corresponds to the covariant form of the Gordon decomposition of the Dirac current (dropping the charge $q$ as a multiplier)
@@ 2.1 Electron’s Wave Function and Space-time Motion in Proper Time
where $\omega_{1}=\omega_{0}/2$, showing that $\phi(0)=A$. In fact, this is just the wave function $\psi_{c}(\tau)$ expressed relative to $X_{c}$ boosted to give the wave function $\psi(x)=\phi(\tau(x))$ relative to $X_{o}$. Deboosting from $X_{o}$ to $X_{c}$ transforms $x$ to $\tau$ because Eq. (2.12) corresponds to the time part of the Lorentz transformation of $x=(ct,\mathbf{x})$ to $x_{c}=(c\tau,\mathbf{r})$, as shown in Beck (2023) in Section 2.2. There is no dependence of the rest-frame wave function on $\mathbf{r}$ because $\mathbf{P}_{c}=\mathbf{0}$.
@@
The 4-vector proper velocity $u^{\mu}(\tau)$ is therefore a superposition of the constant global velocity $v^{\mu}=\pi^{\mu}/m$ and an oscillatory velocity $w^{\mu}(\tau)$ with angular frequency $\omega_{0}$ that is a manifestation of Schrödinger’s zitterbewegung. The same expression for the proper velocity was derived in Barut and Zanghi (1984), Salesi (2002) and Beck (2023) for their respective classical models of the Dirac electron. Notice that with the new notation, the previous velocity fields $\widetilde{u}^{\mu}(x)$ and $\widetilde{w}^{\mu}(x)$ can be written as $u^{\mu}(\tau(x))$ and $w^{\mu}(\tau(x))$; and similarly, $\psi(x)$ is $\phi(\tau(x))$.
@@
where the constant $y^{\mu}(0)=x^{\mu}(0)+\dot{u}^{\mu}(0)/\omega_{0}^{2}=x^{\mu}(0)-z^{\mu}(0)$. Actually, this interpretation is consistent with the Heisenberg operator for rate of change with respect to proper time that was introduced in Beck (1942) and also derived rigorously in Barut and Thacker (1985). Here, the Beck derivative operator gives
@@
and so $\dot{x}^{\mu}(\tau)=\bar{\phi}(\tau)~\hat{\dot{x}}^{\mu}\phi(\tau)=\bar{\phi}(\tau)~\hat{u}^{\mu}\phi(\tau)=u^{\mu}(\tau)$. We also see from its definition in Eq. (2.16) that $z^{\mu}(\tau)$ satisfies the equation of motion $\ddot{z}^{\mu}=-\omega_{0}^{2}z^{\mu}$.
@@ 2.3 Rest-frame Zitter Motion of Electron
We choose origin $C$ to be the point $\mathbf{y}(\tau)$ relative to $X_{o}$ so that it moves with the global motion of the electron relative to $X_{o}$. This means that $C$ is chosen as the point where the momentum of the electron is located ($\mathbf{P}=m\mathbf{v}=m\dot{\mathbf{y}}$, not $m\mathbf{u}$). Clearly, the inertia of the electron is located at the point $C$ and not at the electron’s spatial position $\mathbf{x}(\tau)$ where the charge is located, so $C$ is appropriately called the inertia center (or center of mass) of the electron, although, as we will see, it is also appropriate to call it the spin center. This physical separation of the inertia and charge centers of an electron is a characteristic of the afore-mentioned classical Dirac particle models (see Rivas (2015) for a discussion of this feature). Furthermore, since the free electron is perpetually accelerating (see Eq. (2.15)), it is clear that the electron does not emit or absorb radiation (photons) unless its inertia center accelerates. This opens a new window into the nature of the electron’s inertia and mass that could be further explored.
@@
This is consistent with $\phi^{*}(\tau)\phi(\tau)=\frac{1}{c}\bar{\phi}(\tau)\hat{u}^{0}\phi(\tau)=\frac{1}{c}u^{0}(\tau)=\dot{t}(\tau)=1$ because $t=\tau$ in the rest-frame at $C$. The next equation shows that for a spin state, $|A_{1}|^{2}+|A_{2}|^{2}|=\frac{1}{2}=|A_{3}|^{2}+|A_{4}|^{2}$, so the positive and negative energy solutions are of equal amplitude.
@@
It is readily shown from Eq. (2.29) that $\mathbf{u}\cdot\mathbf{n}=0$ so that the electron’s zitter motion relative to the rest-frame $X_{c}$ lies in a plane orthogonal to the spin direction $\mathbf{n}$. Furthermore, $\mathbf{u}\cdot\mathbf{u}=c^{2}$ so the electron moves at the speed of light, and $\dot{\mathbf{u}}\cdot\dot{\mathbf{u}}=c^{2}\omega_{0}^{2}$, which justify the constraints $C1:u^{\mu}u_{\mu}=0$ and $C2:\dot{u}^{\mu}\dot{u}_{\mu}=-c^{2}\omega_{0}^{2}$ that Beck imposed on his neoclassical model (Beck (2023)). It is shown in Appendix A that $C1$ implies that $\mathbf{r}\cdot\mathbf{r}$ is constant. Thus, the zitter motion is a circle about $C$, which clearly has circular frequency $\omega_{0}$ and therefore radius $r_{0}=c/\omega_{0}=\hbar/(2mc)$.
This circular zitter motion produces a spin $\mathbf{s}$ of magnitude $\hbar/2$ in the direction $\mathbf{n}$ orthogonal to the plane of the motion. Furthermore, if the electron is at space-time point $x$, Eq. (2.20) shows that its angular position in this circular motion is $\omega_{0}\tau(x)=2\theta(x)$, using Eq. (2.12). Thus, as noted in Hestenes (1993), the electron’s angular position in its zitter motion is twice its wave function phase $\theta(x)$; for example, if this phase increases by $\pi$, then the electron will complete a full cycle around its circular zitter path.
@@ 3 Equivalence for Free Electrons of the Dirac Equation and the Classical Dirac Particle Models
This evolution equation is satisfied by the “classical spinor” in the electron model in Barut and Zanghi (1984), which implies that for a free electron, the Dirac wave function written as a function of proper time is a Barut-Zanghi spinor. Beck has proved the converse: given a Barut-Zanghi spinor $\phi(\tau)$, the spinor $\psi(x)=\phi(\tau(x))$, where $\tau(x)$ is defined in Eq. (2.12), satisfies Dirac’s wave equation and so is a Dirac wave function (Beck (2023)). This is obviously true here from reversing the procedure to get Eq. (2.13), that is, $\phi(\tau)$ in this equation is the spinor solution of Eq. (3.1), and so from Eq. (2.12), $\phi(\tau(x))=\psi(x)$, the solution of Dirac‘s equation given in Eq. (2.2).
@@
The free electron models in Hestenes (2010) and in Barut and Zanghi (1984) are also equivalent, as noted in Rodriguez et al. (1993), so the same equivalence with Dirac’s wave equation also applies to Hestenes zitter (light-like) model. Both of these models use a mix of “classical spinors” and classical dynamic variables in their Lagrangian functions. The difference is that Hestenes uses his Space-time Algebra, a Clifford algebra, to express and analyze the Lagrangian function for his zitter model. Note that when there is no em-field, the Pauli dipole term that Hestenes adds to his Lagrangian plays no role in the dynamics.
Barut and Cruz also examine the addition of this dipole term to the Barut-Zanghi Lagrangian function with the goal of empirically producing the “anomalous” magnetic moment that is observed experimentally, which corresponds to a gyromagnetic ratio with $g\approx 2(1+\alpha/(2\pi))$ where $\alpha$ is the usual fine structure constant (Barut and Cruz (1993)). It is important to note, however, that Dirac’s equation does not have an explicit dipole term since it uses only minimal coupling to the em-field but it still gives rise to magnetic and electric dipole moments, albeit with a gyromagnetic ratio corresponding to exactly $g=2$.
@@
giving two coupled second-order equations of motion describing the free electron that are derived from Dirac‘s equation. The equations of motion in Eq. (3.2) are a special case of the neoclassical relativistic mechanics model in Beck (2023) where Newton’s second law in proper time gives the global acceleration $\ddot{y}^{\mu}(\tau)$, consistent with the inertia being located at the inertia center $C$, the point whose spatial position is given by $\mathbf{y}(\tau)$ relative to the inertial reference frame $X_{o}$. His fundamental equations of motion are
@@
where $A=\left(V/c,\mathbf{A}\right)$ is the 4-vector potential for the em-field. This Lagrangian function is just the classical relativistic one with minimal coupling to the em-field (e.g. Goldstein (1959)) except for the additional last term, which turns out to be the contribution from the spin kinetic energy. The same Lagrangian was chosen for a free electron in Riewe (1972) and also in Salesi (2002) for his classical Dirac particle model. We refer to these electron models that are based on this Lagrangian as neoclassical relativistic mechanics models because they use classical relativistic mechanics with the addition of a new spin energy term in the Lagrangian function.
Rivas has presented a similar relativistic model of the electron that uses an inertial observer’s time $t$ rather than the proper time $\tau$ (Rivas (2003)). He derives it from the Frenet-Serret equations for a curve in three-dimensional space. By examining his equations of motion for a free electron relative to the rest-frame $X_{c}$, which is inertial, it can be shown that they are equivalent to the space part of Eq. (3.3) in Beck’s neoclassical model. The time part in Eq. (3.3) is not needed in the rest-frame because observer time $t$ equals proper time $\tau$ there. For the interaction with an em-field, both of these theories use Newton’s second law for the motion of the inertia center, with the Lorentz force evaluated at the charge center. Therefore, the neoclassical model of Beck can be viewed as a covariant form of the relativistic model of Rivas.
Beck introduced a physically-based spin tensor $S$ defined by (Beck (2023))
@@
These same equations were derived in Barut and Zanghi (1984) by applying Dirac’s 4-vector velocity operator and his spin tensor operator to their “classical spinor” $\phi(\tau)$ (they did not give any proof, which is given in Section 4.2 of Beck (2023)). Thus, the simple electron model given in Eq. (3.3) has the same dynamics as the Barut-Zanghi model, and so for a free electron, the same dynamics as Dirac’s electron model. Not only are the equations in Eq. (3.7) implied by Dirac’s equation and operators for a free electron, they also hold for an electron in an em-field provided the dynamic variables $u^{\mu},S^{\mu\nu}$ and $\pi^{\mu}$ are replaced by their corresponding Dirac operators. This can be shown by employing the proper time derivative operator introduced in Beck (1942) to derive the Heisenberg operator equations from Dirac’s equation (e.g. see Barut (1987)).